40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Jun 29 02:33:28 2008, Koji, Update, General, Abs. Len. Meas. ~ Optical setup (I) optical_configuration.pngoptical_layout_ap_table2.pngoptical_layout_ap_table2.pdf
    Reply  Mon Jun 30 05:33:38 2008, Koji, Update, General, Abs. Len. Meas. ~ Optical setup (II) 
       Reply  Tue Jul 1 21:54:05 2008, Koji, Update, General, Re: Abs. Length Meas. setup 
    Reply  Wed Jul 2 15:14:42 2008, steve, Update, General, added beam traps 
Message ID: 619     Entry time: Tue Jul 1 21:54:05 2008     In reply to: 599
Author: Koji 
Type: Update 
Category: General 
Subject: Re: Abs. Length Meas. setup 
I tried to look for the beating in the signal from the PD but I couldn't find. I had the temperature of the laser initially set to 40deg and then slowly increased by one degree. The manual of the laser says the frequency should change by several GHz. The problem is then that our PD is limited to no more than 30Mhz.

Although the two beams seem to overlap quite well, we might still need a better matching of the injected beam.


Alberto


Quote:
o The position of the iris was adjusted so as not to disturub the beam for OMCR CCD.

o The RF spectrum analyzer was returned to the place of the network analyzer.


Quote:

In the process of making this report, I noticed that one of the iris apertures is about disturbing the beam for OMCR CCD. I will check this before I go to Hanford. Also an RF spectrum analyzer is at the AP table. I try to return this near the PSL on Monday morning.
ELOG V3.1.3-